OpenStudy (anonymous):

how would i imput fraction form of a over fraction 1/a?

OpenStudy (anonymous):

a^2

OpenStudy (anonymous):

\[\large\frac{a}{\frac{1}{a}} = \frac{a}{\frac{1}{a}} \cdot \frac{a}{a} = \frac{a^2}{1} = a^2\]

OpenStudy (anonymous):

how do you get the line enbetween the 1 a?

OpenStudy (anonymous):

*[ \frac{a}{ \frac{1}{a} } *] Change * to \

OpenStudy (anonymous):

thank you